OpenStudy (anonymous):

Between 1960 and 1990 the population of a town increased at rate of 0.16% per year. The population P is given by P = 1455 · (1.0016)t where t = 0 corresponds to 1980. Find the population of the town in a. 1960.. So i started off by using the formula above but each time I enter it I get it wrong.

OpenStudy (anonymous):

whats wrong?

OpenStudy (danjs):

t = the time since 1980

OpenStudy (anonymous):

t=20

OpenStudy (danjs):

you want 20 years in the past

OpenStudy (danjs):

moving Left on the x axis , negative

OpenStudy (danjs):

t = -20

OpenStudy (danjs):

But according to that equation, the population is Zero in 1980 then t = 0

OpenStudy (danjs):

sure its not an exponential function?

OpenStudy (anonymous):

is P=1445+1445*(.0016)t

OpenStudy (anonymous):

so it means that in 1980 P was 1445

OpenStudy (danjs):

sure it is not P = 1455*(1.0016)^t

OpenStudy (anonymous):

Yes thats how it is

OpenStudy (danjs):

ok so let t = -20 and find P

OpenStudy (danjs):

\[P = \frac{ 1455 }{ 1.0016^{20} }\]

OpenStudy (anonymous):

I got 1502

OpenStudy (danjs):

i got 1409.21

OpenStudy (danjs):

remember it will be LESS people than in 1980

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ohhhh

OpenStudy (danjs):

1455 / (1.0016)^20

OpenStudy (danjs):

Twenty years in the past, will be a negative exponent, you can move it to the denominatior

OpenStudy (anonymous):

positive 20 I entered a negative So now I got what you got

OpenStudy (anonymous):

I confused is the question 1455*(1.0016)^t???

OpenStudy (danjs):

yes

OpenStudy (anonymous):

:)

OpenStudy (danjs):

Populatioin growth is exponential function

OpenStudy (danjs):

Logistic function really

OpenStudy (anonymous):

So to find 1970 I would use t=10

OpenStudy (danjs):

t = -10

OpenStudy (anonymous):

Which Is 1431

OpenStudy (danjs):

P = 1455(1.0016)^-10 = 1455 / (1.0016)^10

OpenStudy (danjs):

correct

OpenStudy (anonymous):

Oh okay So to find 1990 it would be t=-10

OpenStudy (danjs):

t = +10, but if you move it to the denominator it will be -10

OpenStudy (danjs):

t = years since 1980

OpenStudy (anonymous):

For 1990 t=-10

OpenStudy (anonymous):

and for 1980 it's Just 1455

OpenStudy (danjs):

\[P(1990) = 1455(1.0016)^{+10} = \frac{ 1455 }{ 1.0016^{-10} }\]

OpenStudy (danjs):

Yes for 1980 t = 0, and a number to the zero power is 1

OpenStudy (danjs):

P(1980) = 1455(1)

OpenStudy (anonymous):

This Is fantastic! I got it right and I get the concept!

OpenStudy (danjs):

awesome, if you graphed, 1455(1.0016)^t you will see the growth where t =0 is 1980 it is a curve.

OpenStudy (danjs):

y = 1455(1.0016)^x

OpenStudy (anonymous):

cool

OpenStudy (danjs):

i got to go for a bit, if you want to fan me, ill be back in a little while